# RewardMap: Tackling Sparse Rewards in Fine-grained Visual Reasoning via Multi-Stage Reinforcement Learning

### v1

@@ 6 Conclusion
In Appendix A, we provide a statistical overview of ReasonMap-Plus and its question templates. Appendix B outlines the computation pipeline of the detail reward. Appendix C describes evaluation settings, including difficulty-aware weighting, benchmark datasets, and result sources in Table 2. Finally, Appendix D presents the statement on LLM usage.
@@ A.2 Question Template
Local Counting 1 Please solve the multiple choice problem and put your answer (one of ABCD) in one ” boxed”. According to the subway map, how many intermediate stops are there between stop 1 and stop 2 (except for this two stops)? A) x B) x C) x D) x
Local Counting 2 Please solve the problem and put your answer in one ” boxed”. According to the subway map, how many lines pass through stop 1?
Global Counting Please solve the problem and put your answer in one ” boxed”. According to the subway map, how many subway (metro) lines are there in total?
True or False 1 Please solve the problem and put your answer (only answer yes or no) in one ” boxed”. According to the subway map, is it true that stop 1 is the same line as stop 2?
True or False 2 Please solve the problem and put your answer (only answer yes or no) in one ” boxed”. According to the subway map, is it true that stop 1 is on the line x?
@@ C.2 Details of Evaluation Datasets
    SpatialEval (Wang et al., 2024) targets spatial intelligence across relationships, position, counting, and navigation.
@@
    HRBench (Wang et al., 2025d) evaluates MLLMs on 4K/8K high-resolution images and introduces a training-free enhancement baseline.
@@ C.3 Result Source Summary
We present sources of results in Table 2: $\dagger,\ddagger,\mathdollar$$\S$ correspond to the Seed1.5-VL technical report (Guo et al., 2025b), the Qwen2.5-VL technical report (Bai et al., 2025), the InternVL3.5 technical report (Wang et al., 2025c), the Kimi-VL technical report Team et al. (2025), and the official Hugging Face repository of Qwen2.5-VL-32B-Instruct [^5] https://huggingface.co/Qwen/Qwen2.5-VL-32B-Instruct, respectively.
Appendix D Large Language Model Usage Statement

### v2

@@ 6 Conclusion
Acknowledgement
This paper is supported by Young Scientists Fund of the National Natural Science Foundation of China (NSFC) (No. 62506305), Zhejiang Leading Innovative and Entrepreneur Team Introduction Program (No. 2024R01007), Key Research and Development Program of Zhejiang Province (No. 2025C01026), Scientific Research Project of Westlake University (No. WU2025WF003), Chinese Association for Artificial Intelligence (CAAI) & Ant Group Research Fund - AGI Track (No. 2025CAAI-ANT-13). It is also supported by the research funds of the National Talent Program and Hangzhou Municipal Talent Program.
@@ Acknowledgement
Reproducibility Statement
To ensure reproducibility, we provide the evaluation setup details in Section 5.1 and Appendix C, including hardware and implementation, which facilitates rapid replication.
@@
In Appendix A, we provide a statistical overview of ReasonMap-Plus, question templates, and fine-grained annotation. Appendix B outlines the computation pipeline of the detail reward. Appendix C describes evaluation settings, including difficulty-aware weighting, benchmark datasets, and result sources in Table 2. In Appendix D, we provide more cases to support our claims in Section 5.3. We discuss future work in Appendix E. Finally, Appendix F presents the statement on LLM usage.
@@ A.2 Question Template
Local Counting 1 Please solve the multiple choice problem and put your answer (one of ABCD) in one “ boxed”. According to the subway map, how many intermediate stops are there between stop 1 and stop 2 (except for this two stops)? A) x B) x C) x D) x
Local Counting 2 Please solve the problem and put your answer in one “ boxed”. According to the subway map, how many lines pass through stop 1?
Global Counting Please solve the problem and put your answer in one “ boxed”. According to the subway map, how many subway (metro) lines are there in total?
True or False 1 Please solve the problem and put your answer (only answer yes or no) in one “ boxed”. According to the subway map, is it true that stop 1 is the same line as stop 2?
True or False 2 Please solve the problem and put your answer (only answer yes or no) in one “ boxed”. According to the subway map, is it true that stop 1 is on the line x?
A.3 Fine-Grained Annotation
We provide more details of the process-level annotation in this section. The construction involves two steps. First, we annotate the route–stop structure of each transit map, including the ordered stops for every line, interchange stations, and branching points. These annotations, as listed in the Meta Data, are recorded in a JSON file. An excerpt for Singapore is shown below:
Meta Data “Circle Line”: [“HarbourFront (Transfer)”, “Telok Blangah”, “Labrador Park”, …], “Downtown Line”: […],
Based on the Meta Data, we generate process-level annotations with task-specific Python scripts. For example, in the Local Counting 1 category of ReasonMap-Plus, where the question asks for the number of intermediate stops between two stations, the script identifies the two queried stops, extracts all intermediate stops between them, computes the count as the final answer, and records the intermediate stops as the process-level annotation. Other task types, such as True/False or shortest-path queries, follow a similar procedure.
Example 1 (Local Counting 1) “country”: “singapore”, “city”: “singapore”, “station_1”: “Clarke Quay”, “station_2”: “Farrer Park”, “figure”: “./maps/singapore/singapore.png”, “question”: “According to the map, how many intermediate stops are there between Clarke Quay and Farrer Park (excluding the endpoints)?”, “answer”: “2”, “fine-grained answer”: [“Dhoby Ghaut”, “Little India”], “type”: “counting_1”, “difficulty_city”: “hard”, “city_line_count”: “6”, “city_transfer_count”: “53”, “json”: “./stations/singapore/singapore.json”
Example 2 (True or False 1) “country”: “singapore”, “city”: “singapore”, “station_1”: “Somerset”, “station_2”: “Bras Basah”, “figure”: “./maps/singapore/singapore.png”, “question”: “According to the map, is it true that Somerset is on the same line as Bras Basah?”, “answer”: “No”, “fine-grained answer”: “Somerset”: “North South Line”, “Bras Basah”: “Circle Line” , “type”: “torf_1”, “difficulty_city”: “hard”, “city_line_count”: “6”, “city_transfer_count”: “53”, “json”: “./stations/singapore/singapore.json”
@@ C.2 Details of Evaluation Datasets
    SpatialEval (Wang et al., 2024a) targets spatial intelligence across relationships, position, counting, and navigation.
@@
    HRBench (Wang et al., 2025f) evaluates MLLMs on 4K/8K high-resolution images and introduces a training-free enhancement baseline.
@@ C.3 Result Source Summary
We present sources of results in Table 2: $\dagger,\ddagger,\mathdollar$$\S$ correspond to the Seed1.5-VL technical report (Guo et al., 2025b), the Qwen2.5-VL technical report (Bai et al., 2025b), the InternVL3.5 technical report (Wang et al., 2025e), the Kimi-VL technical report Team et al. (2025), and the official Hugging Face repository of Qwen2.5-VL-32B-Instruct [^5] https://huggingface.co/Qwen/Qwen2.5-VL-32B-Instruct, respectively.
Appendix D Supplementary Results
D.1 Comparison Cases
We provide more comparison cases to support deeper analysis in Section 5.3. In the following case, the prediction from Qwen2.5-VL-7B-Instruct exhibits a visual confusion error at <<wrong1>> and a hallucination at <<wrong2>>, misidentifying three intermediate stops but selecting option A (6). In contrast, the RewardMap-trained model correctly identifies a single intermediate stop and selects the correct option C.
Question: How many intermediate stops are there between A Station and B Station? (<budapest.png> + Format Requirement + {A: 6, B: 8, C: 1, D: 9})
Answer (Qwen2.5-VL-7B-Instruct, Wrong). To determine the number of intermediate stops between A Station and B Station, we need to look at the yellow line on the map, which represents Metro Line 1. 1. Identify the stations on the yellow line between A and B. 2. Count the number of stations, excluding A and B. From the map: – A – Széchenyi Thermal Bath – Bajza Street – B There are 3 intermediate stops between A and B. <<wrong1>> Therefore, the correct answer is: $\boxed{A}$ <<wrong2>>
Answer (Qwen2.5-VL-7B-Instruct + RewardMap, Correct). To solve this problem, we need to count the intermediate stops between A Station and B Station, excluding these two stops. Starting from A: – The next stop is Széchenyi Thermal Bath. Thus, the only intermediate stop between A and B is Széchenyi Thermal Bath. Therefore, the correct answer is: $\boxed{C}$
Appendix E Future Work
Our framework is not limited to transit maps and can be extended to other structured visual domains such as charts and diagrams. Our results support this direction: ChartQA (Masry et al., 2022) improves accuracy from $86.12\%\rightarrow 87.24\%$, Charxiv (Wang et al., 2024b) from $66.41\%\rightarrow 68.90\%$, and InfoVQA (Mathew et al., 2022) from $82.22\%\rightarrow 82.71\%$. These gains suggest that the detail rewards and multi-stage RL scheme generalize to structurally similar tasks. Future work will explore domain-agnostic rewards and broader applications across diverse structured visual reasoning settings.
Efficient methods can further improve the fine-tuned model by our framework (Feng et al., 2024; Bai et al., 2025a; Zhu et al., 2025; Shao et al., 2025a; b; Tao et al., 2025a; b; c). The idea of detail reward can be extended to other fields that require reasoning capability (Zhang et al., 2025d; 2026; Wang et al., 2025d; c; Ai et al., 2025; Wu et al., 2025b; Li et al., 2025e; Wang et al., 2023; Jin et al., 2025; Xiao et al., 2025; Li et al., 2025a; Feng et al., 2026; Wang et al., 2026). Leveraging reinforcement learning for safety alignment (Zhang et al., 2025b; Li & Wang, 2026; Li et al., 2025b) remains a promising direction for further exploration.
Appendix F Large Language Model Usage Statement